Last summer Sam worked for a cycle dealer. The dealer agreed to pay him $210 and a new bike for seven weeks of work. But Sam didn't enjoy the job and quit after four weeks. The dealer gave him $21 and the bike. How much was the bike worth?
(A) $216 (B) $225 (C) $231 (D) $253 (E) None of these

Answer by scott8148(6628)

[(210 + b) / 7] * 4 = 21 + b

840 + 4b = 147 + 7b

693 = 3b